A particle of mass 4.0 kg, initially moving with a velocity of 2.0 m/s collides elastically with a particle of mass 6.0 kg. initially moving with a velocity of -4.0 m/s. What are the velocities of the two particles after the collision?

Grade:12

1 Answers

Raghav Singh
6 years ago
Take right hand side to be positive(+) and left hand side to be negative(-) 
 
Initially, Speed of 4kg block= 2m/s (left to right) and Speed of 6kg block= 4m/s (right to left)
Let speed after collision to be v1 and v2 respectively (left to right both)
 
Conserving momentum we have 
Pinitial = Pfinal
4(2) – 6(4) = 4(v1) + 6(v2) 
-16 = 4(v1) + 6(v2)
2(v1) + 3(v2) = -8 ----equation (i)
 
Now as collison is elastic e=1, i.e
1= velocity of separation after collision/velocity of approach before collision
1= (v2 – v1)/(4+2)
6= v2 – v1 i.e
2(v2) – 2(v1) = 12 ------equation (ii)
 
adding equations (i) and (ii) we have
5(v2) = 4
v2 = 4/5 (as it is +ve it means it is moving from left to right)
Putting value of v2 in equation (ii) we get v1 = -26/5 (as it is -ve it means it is moving from right to left)